# Quillen’s conjecture and unitary groups

### v3

@@ 2.4. Symmetric group and certain subsets of it
We write $G=\operatorname{PGU}\nolimits_{n}(q)$. As we will invoke Theorem 3.9 with $G$ there equal to $G$ here, we check below the hypotheses of such a theorem.
Firstly, as $p\mid(q+1)$, there is an element $\mu$ of order $p$ in $\mathbb{F}^{*}_{q^{2}}$ with $\mu\overline{\mu}=1$. In addition, as $p$ is odd, $p$ divides $q+1$, and $(p,q)\neq(3,2)$, we necessarily have $q\geq 4$. Then we may apply Lemma 4.2 with this value of $\mu$ and obtain $e_{1},e_{2},e_{3},\ldots,e_{n-1}$, $E$, $x$, and $X$, satisfying the conclusions there. As required in Theorem 3.9, $[E]$ is a maximal elementary abelian $p$-subgroup of $G$ by Lemma 4.2(a) and because $\operatorname{m}_{p}(G)=n-1$ by Proposition 2.2. Consider now the following map $h\colon X\to\mathbb{Z}^{*}$,
@@
$^{{x^{\delta}y}\!}{e}_{j_{k}}=z_{k}\cdot e_{n-1}^{m^{n-1}_{k}}\cdots e_{k}^{m^{k}_{k}},$
for some integers $m^{i}_{k}$ and some elements $z_{k}\in Z$. Conjugating by $x^{-\delta}$, and using, if $\delta=1$, Lemma 4.2(d) and Equation (13), we obtain that
(15) $^{{y}\!}{e}_{j_{k}}=z_{k}\cdot e_{n-1}^{m^{n-1}_{k}}\cdots e_{k}^{m^{k}_{k}}.$
@@
Note that the base case is vacuous and assume that Equation (16) holds for $k$ with $n\geq k\geq 3$. Then Equation (10) of Proposition 2.8, with the $k$ there equal to the $k$ here, shows that $2\leq y(k)\leq k$. On the other hand, we have, by hypothesis, that $j_{k-1}\leq k-1$. Then, checking against Equation (15), with the $k$ there equal to k1k-1 here, we find that the only solution is $y(k)=k$ and $j_{k-1}={k-1}$. Hence, Equation (16) holds for k1k-1.
@@
$^{{x^{\delta}}\!}{e}_{1}=z_{1}\cdot e_{n-1}^{m^{n-1}_{1}}\cdots e_{2}^{m^{2}_{1}}e_{1}^{m^{1}_{1}}$
for some integers $m^{i}_{1}$ and some $z_{1}\in Z$. If $\delta=1$, we get a contradiction with Lemma 4.2(d) and Equation (13). Thus, $\delta=0$ and we are done.
@@
and $D^{-}([x^{\delta}y],\mathbf{i},\mathbf{j})$ and $D^{+}([x^{\delta}y],\mathbf{i},\mathbf{j})$ as the corresponding sums. Note that, by the first part of Proposition 2.8(a),
(17) $D([x^{\delta}y],\mathbf{i},\mathbf{j})=D^{-}([x^{\delta}y],\mathbf{i},\mathbf{j})+D^{+}([x^{\delta}y],\mathbf{i},\mathbf{j}),$
where $D([x^{\delta}y],\mathbf{i},\mathbf{j})$ was defined in the proof of Lemma 3.11, and we will prove here that $D^{-}([x^{\delta}y],\mathbf{i},\mathbf{j})=D^{+}([x^{\delta}y],\mathbf{i},\mathbf{j})=0$.
For the case of $D^{-}([x^{\delta}y],\mathbf{i},\mathbf{j})$, we note that the element $y_{j_{1}}\in Y$ belongs to $C_{\operatorname{GU}\nolimits_{n}}(q)(E_{j_{1}})=C_{\operatorname{GU}\nolimits_{n}(q)}(\tau_{\mathbf{j}})$ by Lemma 4.2(b), so that $D^{-}([x^{\delta}y],\mathbf{i},\mathbf{j})=0$ follows as in the proof of Lemma 3.11(c) employing Equation (14).
To show that $D^{+}([x^{\delta}y],\mathbf{i},\mathbf{j})=0$, we consider the map
@@
To see that $\Omega$ is well defined, note that, if $x^{\delta^{\prime}}y^{\prime}]\in\mathcal{D}^{+}([x^{\delta}y],\mathbf{i},\mathbf{j})$, then $y^{\prime}y_{j_{1}}\in Y^{+}$, $y^{\prime}\in\partial Y^{-}$ by Definition 2.6, and $y(\{j_{1},j_{1}+1\})=\{1,2\}$ by Proposition 2.8(b). Thus, by Equations (13) and (12), we obtain that the first two diagonal entries of $^{{y^{\prime}}\!}{e}_{i}$ are equal for $i\neq j_{1}$. Thus, by Lemma 4.2(d), $x\in C_{\operatorname{GU}\nolimits_{n}(q)}({}^{{y^{\prime}}\!}{E_{j_{1}}})=C_{\operatorname{GU}\nolimits_{n}(q)}({}^{{y^{\prime}}\!}{\tau_{\mathbf{j}}})$ and
@@
so that $\Omega$ is well defined. Moreover, the map $\Omega$ is easily seen to be a non-trivial involution by Lemma 4.2(c). Thus, we may partition $\mathcal{D}^{+}([x^{\delta}y],\mathbf{i},\mathbf{j})$ into pairwise disjoint subsets of the form $\{[x^{\delta^{\prime}}y^{\prime}],[x^{1-\delta^{\prime}}y^{\prime}]\}$ and, in addition, we have
$h]([x^{\delta^{\prime}}y^{\prime}])+[h]([x^{1-\delta^{\prime}}y^{\prime}])=h(x^{\delta^{\prime}}y^{\prime})+h(x^{1-\delta^{\prime}}y^{\prime})=\operatorname{sgn}(y^{\prime})((-1)^{\delta^{\prime}}+(-1)^{1-\delta^{\prime}})=0$
by Equation (14). We are done. ∎
@@
$\begin{cases}\lambda^{q+1}=1\text{, and}\\ \operatorname{ord}(\lambda)=\operatorname{ord}(\Lambda)=r\text{ for a prime $r\neq 2,p$.}\end{cases}$
@@
By Fermat’s little theorem, and because $p|(q+1)=(s^{pl}+1)$, we have that $p|(s^{l}+1)$. Now, by Zsigmondy’s Theorem, there exists a prime $r$ that divides $s^{2pl}-1$ and does not divide $s^{kl}-1$ for $1\leq k<2p$. This works but for $2p=1,2$ or $(p,q)=(3,8)$, and these cases are excluded by assumption. As $1\leq p<2p$, we get that $r$ does not divide $s^{pl}-1$ and, as $s^{2pl}-1=(s^{pl}-1)(s^{pl}+1)$, we deduce that $r$ divides $s^{pl}+1$. In particular, $r\neq 2=(s^{pl}+1)-(s^{pl}-1)$. As $1\leq 2<2p$, $r$ does not divide $s^{2l}-1$ and, as $p\mid(s^{l}+1)$ and $(s^{l}+1)\mid(s^{2l}-1)$, we deduce that $r\neq p$. To finish, define $\lambda$ as any element of order $r$ of $\mathbb{F}_{q^{2}}$. ∎
The next result serves as preparation for Theorem 4.7, and the only conclusion below that fails for $n=2$ is the third part of point (e). In fact, the case of $\operatorname{PGU}\nolimits_{2}(q)$ extended by field automorphisms of order $p$ requires slightly different techniques, and we have omitted this case from the present work. In point (b) below, we recall that $E^{\Phi}_{i}$ is the subgroup of $E^{\Phi}$ spanned by the $e_{k}$ with $k\neq i$.
@@
If $n\geq 3$, $q=s^{pl}$, $p$ is odd, $p\mid(q+1)$, and $(p,q)\neq(3,8)$, then there exist elements of order $p$, $e_{1},e_{2},e_{3},\ldots,e_{n-1},e_{n}=\Phi$, in $\operatorname{GU}\nolimits_{n}(q)\langle\Phi\rangle$, such that, for $E^{\Phi}=\langle e_{1},e_{2},\ldots,e_{n-1},e_{n}\rangle$, we have,
@@
    $E^{\Phi}]$ is an elementary abelian $p$-subgroup of $\operatorname{PGU}\nolimits_{n}(q)\langle\Phi\rangle$ of rank $n$, and
@@
    $y_{i}\in C_{\operatorname{GU}\nolimits_{n}(q)}(E^{\Phi}_{i})$ for $1\leq i\leq n-1$.
Moreover, there exists a non-diagonal element $x\in\operatorname{GU}\nolimits_{n}(q)$ and a diagonal element $d\in GU_{n}(q)$ such that, if we set $X=\{x^{\delta}y\mid\delta=0,1,y\in Y^{-}\}$, we have
@@
If we set $E^{\Phi}=\langle e_{1},\ldots,e_{n-1},e_{n}\rangle$, where $e_{n}=\Phi$, then it is clear that $E^{\Phi}]$ is an elementary abelian $p$-subgroup of $\operatorname{PGU}\nolimits_{n}(q)\langle\Phi\rangle$ of rank $n$ as $E^{\Phi}\cap Z=E\cap Z$ and $E\cap Z=1$ by Lemma 4.2(a). Thus, point (a) holds. Points (c) and (d) are consequence of points (c) and (d) in Lemma 4.2, respectively. The first and second parts of point (e) are exactly Equation (18), and the fourth part of point (e) follows from that $y_{i},\Phi]=1$ by Definition 4.1. Then the fourth part of point of (e) and Lemma 4.2(b) show that point (b) holds.
Now, by Lemma 4.5, we obtain $\lambda$ and $\Lambda$ satisfying the properties there, and we define $d=\operatorname{diag}(1,\ldots,1,\lambda)$. Then the third part of point (e) is consequence of Equation (6) and the fact that $n\geq 3$, and the fifth part of point (e) is a consequence of that $d$ is diagonal and of Equation (12). Thus, we are left with proving the last part of point (e) and, to that aim, we perform the following computation,
$^{{d}\!}{\Phi}=d\Phi(d^{-1})\Phi=\operatorname{diag}(1,\ldots,1,\Lambda)\Phi.$
Assume now that $^{{d}\!}{\Phi}=\operatorname{diag}(1,\ldots,1,\Lambda)\Phi=ze\Phi^{m}$, where $m$ is an integer, $e\in E$, and $z\in Z$. Then we immediately deduce that $m=1$ and, after raising to the $p$-th power, that $\operatorname{diag}(1,\ldots,1,\Lambda^{p})=z^{p}$. As $n\geq 3$, it is clear that $z^{p}=1$, and hence that $\Lambda^{p}=1$, but the order of $\Lambda$ is a prime not equal to $p$. ∎
@@
We are assuming that $\operatorname{PGU}\nolimits_{n}(q)$ admits a field automorphisms of odd order $p$, so that we may write $q=s^{pl}$. As two order-$p$ subgroups of field automorphisms of $\operatorname{PGU}\nolimits_{n}(q)$ are $\operatorname{PGU}\nolimits_{n}(q)$-conjugate by Subsection 2.1, we may assume without loss of generality that $G=\operatorname{PGU}\nolimits_{n}(q)\langle\Phi\rangle$.
Now, we apply Lemma 4.6 and obtain $e_{1},e_{2},e_{3},\ldots,e_{n-1},e_{n}$, $E^{\Phi}$, $x$, $d$, and $X$ satisfying the conclusions there. Next, consider the map $h\colon X\to\mathbb{Z}^{*}$ given by Equation (14) and the (n1)(n-1)-chain provided by Definition 3.8 for the group $G$,
$\mathbf{C}=\mathbf{C}_{[E^{\Phi}],[X],[h]},$
as well as the chain obtained by conjugating $\mathbf{C}$ by $d$,
$\mathbf{C}^{\prime}={}^{{d}\!}{\mathbf{C}}=\mathbf{C}_{[{}^{{d}\!}{E}^{\Phi}],[{}^{{d}\!}{X}],[{}^{{d}\!}{h}]},$
@@
$^{{d}\!}{E}^{\Phi}={}^{{d}\!}{\langle}e_{1},e_{2},e_{3},\ldots,e_{n-1},e_{n}\rangle=\langle e_{1},e_{2},e_{3},\ldots,e_{n-1},{}^{{d}\!}{\Phi}\rangle$
@@
We proof below that $\mathbf{C}-\mathbf{C}^{\prime}$ is is a non-trivial cycle, showing that the conclusions of the statement are valid. For $\mathbf{i}\in\mathcal{T}^{n}$, we use the notations $\sigma_{\mathbf{i}}$, $\tau_{\mathbf{i}}$, $C_{\cdot,\mathbf{i}}$, $\mathcal{C}(\cdot,\mathbf{i})$, etc, of Section 3 for $\mathbf{C}$, and their primed versions $\sigma^{\prime}_{\mathbf{i}}$, $\tau^{\prime}_{\mathbf{i}}$, $C^{\prime}_{\cdot,\mathbf{i}}$, $\mathcal{C}^{\prime}(\cdot,\mathbf{i})$, etc, for $\mathbf{C}^{\prime}$. Note that $E^{\Phi}]$ is a maximal elementary abelian $p$-subgroup of $G$ by Lemma 4.6(a) and because $\operatorname{m}_{p}(G)=n$ by Proposition 2.2. In particular, $\mathbf{C}-\mathbf{C}^{\prime}$ cannot be a boundary.
@@
$\langle[{}^{{d^{\epsilon}}\!}{e}_{j_{n-1}}]\rangle<\langle[{}^{{d^{\epsilon}}\!}{e}_{j_{n-1}}],[{}^{{d^{\epsilon}}\!}{e}_{j_{n-2}}]\rangle<\cdots<\langle[{}^{{d^{\epsilon}}\!}{e}_{j_{n-1}}],[{}^{{d^{\epsilon}}\!}{e}_{j_{n-2}}],\ldots,[{}^{{d^{\epsilon}}\!}{e}_{j_{1}}]\rangle<[{}^{{d^{\epsilon}}\!}{E}^{\Phi}],$
@@
$\sigma_{\mathbf{i}}]=\langle[e_{n-1}]\rangle<\langle[e_{n-1}],[e_{n-2}]\rangle<\cdots<\langle[e_{n-1}],[e_{n-2}],\ldots,[e_{1}]\rangle<[E^{\Phi}].$
@@
The (n1)(n-1)-chain $\mathbf{C}-\mathbf{C}^{\prime}$ is a cycle. For tuples $\mathbf{i},\mathbf{j}\in\mathcal{T}^{n}$, where $\mathbf{j}=[j_{1},\ldots,j_{n-1}]\in\mathcal{T}^{n}$, and $x^{\delta}y\in X$, we define, as in the proof of Lemma 3.11,
$\mathcal{D}([x^{\delta}y],\mathbf{i},\mathbf{j})$ $=\{[x^{\delta^{\prime}}y^{\prime}]\in[X]~|~[{}^{{x^{\delta^{\prime}}y^{\prime}}\!}{\tau}_{\mathbf{j}}]=[{}^{{x^{\delta}y}\!}{\tau}_{\mathbf{i}}]\},$
$\mathcal{D}^{\prime}([x^{\delta}\ {}^{{d}\!}{y}],\mathbf{i},\mathbf{j})$ $=\{[x^{\delta^{\prime}}\ {}^{{d}\!}{y^{\prime}}]\in[{}^{{d}\!}{X}]~|~[{}^{{x^{\delta^{\prime}}\ {}^{{d}\!}{y^{\prime}}}\!}{\tau}^{\prime}_{\mathbf{j}}]=[{}^{{x^{\delta}\ {}^{{d}\!}{y}}\!}{\tau}^{\prime}_{\mathbf{i}}]\},$
and we write $D([x^{\delta}y],\mathbf{i},\mathbf{j})$ and $D^{\prime}([x^{\delta}\ {}^{{d}\!}{y}],\mathbf{i},\mathbf{j})$ for the corresponding sums.
If $j_{1}=1,\ldots,n-1$, we may define, as we did in the proof of Theorem 4.3, the sums $D^{-}([x^{\delta}y],\mathbf{i},\mathbf{j})$ and $D^{+}([x^{\delta}y],\mathbf{i},\mathbf{j})$. Then the arguments in the proof of Theorem 4.3 show that $D^{-}([x^{\delta}y],\mathbf{i},\mathbf{j})=0$, where we use Lemma 4.6(b) instead of Lemma 4.2(b), and that $D^{+}([x^{\delta}y],\mathbf{i},\mathbf{j})=0$, where we need to employ that, by the second part of Lemma 4.6(e), $x,\Phi]=1$. Thus, by the analogue here of Equation (17), we have that $D([x^{\delta}y],\mathbf{i},\mathbf{j})=0$. Finally, note that $\mathcal{D}^{\prime}([x^{\delta}\ {}^{{d}\!}{y}],\mathbf{i},\mathbf{j})={}^{{[d]}\!}{\mathcal{D}}([x^{\delta}y],\mathbf{i},\mathbf{j})$, and thus we also have $D^{\prime}([x^{\delta}\ {}^{{d}\!}{y}],\mathbf{i},\mathbf{j})=0$ for $j_{1}=1,\ldots,n-1$.
If $j_{1}=n$, we show that $D([x^{\delta}y],\mathbf{i},\mathbf{j})+D^{\prime}([x^{\delta}\ {}^{{d}\!}{y}],\mathbf{i},\mathbf{j})=0$. In fact, consider the map
@@
$\langle[e_{n-1}]\rangle<\langle[e_{n-1}],[e_{n-2}]\rangle<\cdots<\langle[e_{n-1}],\ldots,[e_{2}]\rangle<\langle[e_{n-1}],\ldots,[e_{2}],[e_{1}]\rangle<[E^{\Phi}],$
@@
We dedicate this section to provide details for the proof of Theorem B stated in the introduction, and we employ [3, Lemma 4.7, Proposition 4.8].

### v5

@@ 2.4. Symmetric group and certain subsets of it
We check the hypotheses of Theorem 3.9 for $G=\operatorname{PGU}\nolimits_{n}(q)$. Firstly, as $p\mid(q+1)$, there is an element $\mu$ of order $p$ in $\mathbb{F}^{*}_{q^{2}}$ with $\mu\overline{\mu}=1$. In addition, as $p$ is odd, $p$ divides $q+1$, and $(p,q)\neq(3,2)$, we necessarily have $q\geq 4$. Then we may apply Lemma 4.2 with this value of $\mu$ and obtain $e_{1},e_{2},e_{3},\ldots,e_{n-1}$, $E$, $x$, and $X$, satisfying the conclusions there. As required in Theorem 3.9, $[E]$ is a maximal elementary abelian $p$-subgroup of $G$ by Lemma 4.2(a) and because $\operatorname{m}_{p}(G)=n-1$ by Proposition 2.2. Consider now the following map $h\colon X\to\mathbb{Z}^{*}$,
@@
$^{{x^{\delta}y}\!}{e}_{j_{k}}=z_{k}\cdot e_{n-1}^{m_{k,n-1}}\cdots e_{k}^{m_{k,k}},$
for some integers $m_{k,i}$ and some elements $z_{k}\in Z$. Conjugating by $x^{-\delta}$, and using, if $\delta=1$, Lemma 4.2(d) and Equation (12), we obtain that
(15) $^{{y}\!}{e}_{j_{k}}=z_{k}\cdot e_{n-1}^{m_{k,n-1}}\cdots e_{k}^{m_{k,k}}.$
@@
We note that the base case $k=n$ is vacuous and assume that Equation (16) holds for $k$ with $n\geq k\geq 3$. Then Equation (8) of Proposition 2.8, with the $k$ there equal to the $k$ here, shows that $2\leq y(k)\leq k$. On the other hand, we have, by hypothesis, that $j_{k-1}\leq k-1$. Then, checking against Equation (15), with the $k$ there equal to k1k-1 here, we find that the only solution is $y(k)=k$ and $j_{k-1}={k-1}$. Hence, Equation (16) holds for k1k-1.
@@
$^{{x^{\delta}}\!}{e}_{1}=z_{1}\cdot e_{n-1}^{m_{1,n-1}}\cdots e_{2}^{m_{1,2}}e_{1}^{m_{1,1}}$
for some integers $m_{1,i}$ and some $z_{1}\in Z$. If $\delta=1$, we get a contradiction with Lemma 4.2(d) and Equation (12). Thus, $\delta=0$ and we are done.
@@
and denote by $D^{-}([x^{\delta}y],\mathbf{i},\mathbf{j})$ and $D^{+}([x^{\delta}y],\mathbf{i},\mathbf{j})$ the corresponding sums. Then, by the first and second parts of Proposition 2.8(a), we obtain that, see Equation (10),
$D([x^{\delta}y],\mathbf{i},\mathbf{j})=D^{-}([x^{\delta}y],\mathbf{i},\mathbf{j})+D^{+}([x^{\delta}y],\mathbf{i},\mathbf{j}),$
and, by Lemma 3.12(c), it is enough to show that both summands above are zero.
Now, we obtain that $D^{-}([x^{\delta}y],\mathbf{i},\mathbf{j})=D^{*}([x^{\delta}y],\mathbf{i},\mathbf{j})$ from Equation (13), and that $D^{*}([x^{\delta}y],\mathbf{i},\mathbf{j})=0$ from Lemma 4.2(b), Equation (14), and Lemma 3.12(b,b2). To show that $D^{+}([x^{\delta}y],\mathbf{i},\mathbf{j})=0$, we consider the map
@@
To see that $\Omega$ is well defined, note that, if $x^{\delta^{\prime}}y^{\prime}]\in\mathcal{D}^{+}([x^{\delta}y],\mathbf{i},\mathbf{j})$, then $y^{\prime}y_{j_{1}}\in Y^{+}$, $y^{\prime}\in\partial Y^{-}$ by Definition 2.6, and $y^{\prime}(\{j_{1},j_{1}+1\})=\{1,2\}$ by Proposition 2.8(b). Thus, by Equations (12) and (11), we obtain that the first two diagonal entries of $^{{y^{\prime}}\!}{e}_{i}$ are equal for $i\neq j_{1}$. Thus, by Lemma 4.2(d), $x\in C_{\operatorname{GU}\nolimits_{n}(q)}({}^{{y^{\prime}}\!}{E_{j_{1}}})=C_{\operatorname{GU}\nolimits_{n}(q)}({}^{{y^{\prime}}\!}{\tau_{\mathbf{j}}})$ and
@@
so that $\Omega$ is well defined. Moreover, the map $\Omega$ is a non-trivial involution by Lemma 4.2(c), and thus we are done by Lemma 3.11 and Equation (14). ∎
@@
$\begin{cases}\lambda^{q+1}=1\text{, and}\\ \operatorname{ord}(\lambda)=\operatorname{ord}(\Lambda)=r\text{ for a prime $r\neq 2,3,5,p$.}\end{cases}$
@@
By Fermat’s little theorem, and because $p|(q+1)=(s^{pl}+1)$, we have that $p|(s^{l}+1)$. Now, by Zsigmondy’s Theorem, there exists a prime $r$ that divides $s^{2pl}-1$ and does not divide $s^{kl}-1$ for $1\leq k<2p$. This works but for $2p=1,2$ or $(p,q)=(3,8)$, and these cases are excluded by assumption. As $1\leq p<2p$, we get that $r$ does not divide $s^{pl}-1$ and, as $s^{2pl}-1=(s^{pl}-1)(s^{pl}+1)$, we deduce that $r$ divides $s^{pl}+1$. As $1\leq 2<2p$, $r$ does not divide $s^{2l}-1$ and, as $p\mid(s^{l}+1)$ and $(s^{l}+1)\mid(s^{2l}-1)$, we deduce that $r\neq p$. In addition, we have, by construction, that the order of $s^{l}$ modulo $r$ is $2p$. Then, by Fermat’s little theorem again, we deduce that $2p\mid(r-1)$, so that $r\geq 1+2p\geq 7$ and $r\neq 2,3,5$. To finish, define $\lambda$ as any element of order $r$ of $\mathbb{F}_{q^{2}}$. ∎
The next result serves as preparation for Theorem 4.7, and the only conclusion below that fails for $n=2$ is the third part of point (e). In fact, the case of $\operatorname{PGU}\nolimits_{2}(q)$ extended by field automorphisms of order $p$ requires slightly different techniques, and we have omitted this case from the present work. In point (b) below, we recall that $(E_{\Phi})_{i}$ is the subgroup of $E_{\Phi}$ spanned by the $e_{k}$ with $k\neq i$.
@@
If $n\geq 3$, $q=s^{pl}$, $p$ is odd, $p\mid(q+1)$, and $(p,q)\neq(3,8)$, then there exist elements of order $p$, $e_{1},e_{2},e_{3},\ldots,e_{n-1},e_{n}=\Phi$, in $\operatorname{GU}\nolimits_{n}(q)\langle\Phi\rangle$, such that, for $E_{\Phi}=\langle e_{1},e_{2},\ldots,e_{n-1},e_{n}\rangle$, we have,
@@
    $E_{\Phi}]$ is an elementary abelian $p$-subgroup of $\operatorname{PGU}\nolimits_{n}(q)\langle\Phi\rangle$ of rank $n$, and
@@
    $y_{i}\in C_{\operatorname{GU}\nolimits_{n}(q)}((E_{\Phi})_{i})$ for $1\leq i\leq n-1$.
Moreover, there exists an element $x\in\operatorname{GU}\nolimits_{n}(q)$ and a diagonal element $d\in\operatorname{SU}\nolimits_{n}(q)$ such that, if we set $X=\{x^{\delta}y\mid\delta=0,1,y\in Y^{-}\}$, we have
@@
If we set $E_{\Phi}=\langle e_{1},\ldots,e_{n-1},e_{n}\rangle$, where $e_{n}=\Phi$, then it is clear that $E_{\Phi}]$ is an elementary abelian $p$-subgroup of $\operatorname{PGU}\nolimits_{n}(q)\langle\Phi\rangle$ of rank $n$, as $E_{\Phi}\cap Z=E\cap Z$ and $E\cap Z=1$ by Lemma 4.2(a), where we recall that $Z=Z(\operatorname{GU}\nolimits_{n}(q))$. Thus, point (a) holds. Points (c) and (d) are consequence of points (c) and (d) in Lemma 4.2, respectively. The first and second parts of point (e) are exactly Equation (17), and the fourth part of point (e) follows from that $y_{i},\Phi]=1$ by Definition 4.1. Then the fourth part of point of (e) and Lemma 4.2(b) show that point (b) holds.
Now, from Lemma 4.5, we obtain $\lambda$ and $\Lambda$ satisfying the properties there, and we define
$d=\begin{cases}\operatorname{diag}(\lambda,\lambda,\lambda^{-2})&n=3,\\ \operatorname{diag}(1,\ldots,1,\lambda,\lambda^{-1})&n\geq 4,\end{cases}$
so that $d\in\operatorname{SU}\nolimits_{n}(q)$. Then the third part of point (e) is consequence of Equation (4), and the fifth part of point (e) is a consequence of that $d$ is diagonal and of Equation (11). Thus, we are left with proving the last part of point (e) and, to that aim, we perform the following computation,
$^{{d}\!}{\Phi}=d\Phi(d^{-1})\Phi=\begin{cases}\operatorname{diag}(\Lambda,\Lambda,\Lambda^{-2})\Phi&n=3,\\ \operatorname{diag}(1,\ldots,1,\Lambda,\Lambda^{-1})\Phi&n\geq 4.\end{cases}$
Assume now that $^{{d}\!}{\Phi}=ze\Phi^{m}$, where $0\leq m\leq p-1$, $e\in E$, and $z\in Z$. Then we immediately deduce that $m=1$ and, after raising to the $p$-th power, that $\Lambda^{p}=1$ if $n\geq 4$ or that $\Lambda^{3p}=1$ if $n=3$. But the order of $\Lambda$ is a prime not equal to $p$ or $3$ and we are done. ∎
@@
We are assuming that $\operatorname{PGU}\nolimits_{n}(q)$ admits a field automorphisms of odd order $p$, so that we may write $q=s^{pl}$. As two order-$p$ subgroups of field automorphisms of $\mathrm{Aut}(\operatorname{PSU}\nolimits_{n}(q))$ are $\operatorname{PGU}\nolimits_{n}(q)$-conjugate by Subsection 2.1, we may assume without loss of generality that $G=\operatorname{PGU}\nolimits_{n}(q)\langle\Phi\rangle$.
Now, we apply Lemma 4.6 and obtain $e_{1},e_{2},e_{3},\ldots,e_{n-1},e_{n}$, $E_{\Phi}$, $x$, $d$, and $X$ satisfying the conclusions there. Next, consider the map $h\colon X\to\mathbb{Z}^{*}$ given by Equation (14) and the (n1)(n-1)-chain provided by Definition 3.8 for the group $G$,
$\mathbf{C}=\mathbf{C}_{[E_{\Phi}],[X],[h]},$
as well as the chain obtained by conjugating $\mathbf{C}$ by $[d]$,
$\mathbf{C}^{\prime}={}^{{[d]}\!}{\mathbf{C}}=\mathbf{C}_{[{}^{{d}\!}{E}_{\Phi}],[{}^{{d}\!}{X}],[{}^{{d}\!}{h}]},$
@@
$^{{d}\!}{E}_{\Phi}={}^{{d}\!}{\langle}e_{1},e_{2},e_{3},\ldots,e_{n-1},e_{n}\rangle=\langle e_{1},e_{2},e_{3},\ldots,e_{n-1},{}^{{d}\!}{\Phi}\rangle$
@@
We proof below that $\mathbf{C}-\mathbf{C}^{\prime}$ is is a non-trivial cycle, showing that the conclusions of the statement are valid. For $\mathbf{i}\in\mathcal{T}^{n}$, we use the notations $\sigma_{\mathbf{i}}$, $\tau_{\mathbf{i}}$, $\mathcal{C}(\cdot,\mathbf{i})$, $\mathcal{D}(\cdot,\mathbf{i},\mathbf{j})$, etc, of Section 3 and Equation (10) for $\mathbf{C}$, and their primed versions $\sigma^{\prime}_{\mathbf{i}}$, $\tau^{\prime}_{\mathbf{i}}$, $\mathcal{D}^{\prime}(\cdot,\mathbf{i},\mathbf{j})$, etc, for $\mathbf{C}^{\prime}$. Note that $E_{\Phi}]$ is a maximal elementary abelian $p$-subgroup of $G$ by Lemma 4.6(a) and because $\operatorname{m}_{p}(G)=n$ by Proposition 2.2. In particular, $\mathbf{C}-\mathbf{C}^{\prime}$ cannot be a boundary.
@@
$\langle[{}^{{d^{\epsilon}}\!}{e}_{j_{n-1}}]\rangle<\langle[{}^{{d^{\epsilon}}\!}{e}_{j_{n-1}}],[{}^{{d^{\epsilon}}\!}{e}_{j_{n-2}}]\rangle<\cdots<\langle[{}^{{d^{\epsilon}}\!}{e}_{j_{n-1}}],[{}^{{d^{\epsilon}}\!}{e}_{j_{n-2}}],\ldots,[{}^{{d^{\epsilon}}\!}{e}_{j_{1}}]\rangle<[{}^{{d^{\epsilon}}\!}{E}_{\Phi}],$
@@
$\sigma_{\mathbf{i}}]=\langle[e_{n-1}]\rangle<\langle[e_{n-1}],[e_{n-2}]\rangle<\cdots<\langle[e_{n-1}],[e_{n-2}],\ldots,[e_{1}]\rangle<[E_{\Phi}].$
@@
The (n1)(n-1)-chain $\mathbf{C}-\mathbf{C}^{\prime}$ is a cycle. Consider tuples $\mathbf{i},\mathbf{j}\in\mathcal{T}^{n}$, where $\mathbf{j}=[j_{1},\ldots,j_{n-1}]\in\mathcal{T}^{n}$, and an element $x^{\delta}y\in X$.
If $j_{1}=1,\ldots,n-1$, then the arguments in the proof of Theorem 4.3 show that $D([x^{\delta}y],\mathbf{i},\mathbf{j})=0$, where we use Lemma 4.6(b) instead of Lemma 4.2(b) and that, by the second part of Lemma 4.6(e), $x,\Phi]=1$. Moreover, as $\mathcal{D}^{\prime}([x^{\delta}\ {}^{{d}\!}{y}],\mathbf{i},\mathbf{j})$ is the $[d]$-conjugate of $\mathcal{D}([x^{\delta}y],\mathbf{i},\mathbf{j})$, we deduce that $D^{\prime}([x^{\delta}\ {}^{{d}\!}{y}],\mathbf{i},\mathbf{j})=0$ too.
If $j_{1}=n$, we show that $D([x^{\delta}y],\mathbf{i},\mathbf{j})+D^{\prime}([x^{\delta}\ {}^{{d}\!}{y}],\mathbf{i},\mathbf{j})=0$ by means of the map
@@
$\langle[e_{n-1}]\rangle<\langle[e_{n-1}],[e_{n-2}]\rangle<\cdots<\langle[e_{n-1}],\ldots,[e_{2}]\rangle<\langle[e_{n-1}],\ldots,[e_{2}],[e_{1}]\rangle<[E_{\Phi}],$
@@
We dedicate this section to provide details for the proof of Theorem B stated in the introduction, and we employ [4, Proposition 4.8].